There was a power blip around 11:50am that affected the 3rd floor machine room. The power blip took down the CS domain controllers and print servers. All services have been restored now.
Two GPFS NFS servers failed simultaneously at around 1pm today, causing file services to be unavailable for about 40 minutes. The servers, crows and runts, serve different cluster groups; crows serves the grid and runts serves the internal department network. Runts' failure was a kernel lock-up which prevented the normal automatic failover behavior. The technical ...
The Computer Science Department event calendar is now hosted on Brown's campus-wide event calendar system. The "topical area" of Computer Science is where you will find our events both past and future. The CS website still displays upcoming events on the home page and on the events page using an ical feed from university ...
One of our GPFS NFS servers, peeps, hung at around 10:20am today. File services were unresponsive for about 30 minutes. We are investigating the cause. All services should now be back to normal.
Tstaff will decommission the CCV Backup Service of the GPFS filesystem on the end of business day, Friday, April, 28, 2017. We will no longer have access to the CCV tape library after that date. The result is we will not be able to fulfill data restore requests for data older than 6 months. Please ...
The certificate for the filemaker server, filemakera.ad.brown.edu, is about to expire. In order to perform maintenance on the server, all databases on the filemakera.ad.brown.edu have been migrated to filemakerh.ad.brown.edu. The syschange is scheduled for the morning of Thursday, March 9th. After Thursday, March 9th, the CS ...
There was brief departament cluster NFS blip at 12pm today. Two of the dept cluster nfs servers, smarties and peeps, rebooted themselves causing the nfs blip. We are investigating the root cuase of the reboot.
There was a broken water pipe on the 3rd floor near the network closet. One of the network switch, cit-cs-as.net.brown, was affected and is offlined. The following CIT rooms are affected this switch outage: 115, 121, 132, and 134.
We are investigating further and will update this post as we learn more.
There was a short network outage at 5pm today, affecting some of our user managed machines in Room 310 and the department's Print Host. While disabling a set of ports on the switch a port we were not suppose to have access to was also disabled, which severed that switch's communications with the ...
The version of Chromium currently installed on our managed machines is affected by the bug described at the link above. This bug prevents the browser from completing the necessary steps for a secure connection, which results in a generic error message.
For those using Chromium, until ...
- January 2017
- February 2017
- March 2017
- April 2017
- May 2017
- June 2017
- July 2017
- August 2017
- September 2017
- October 2017
- November 2017
- December 2017